# Defining subsystems in Hilbert spaces with non-Euclidean metric

### v2

@@ II.2 Partitions of a system
Let $\mathcal{A}_{\pi}^{1}=\{A\in\mathcal{A}:\eta\pi(A)\eta^{-1}=\mathcal{O}_{A}\otimes\mathbb{1}\}$ and $\mathcal{A}_{\pi^{\prime}}^{1}=\{A\in\mathcal{A}:\eta^{\prime}\pi_{G^{\prime}}(A)\eta^{\prime-1}=\mathcal{O}_{A}^{\prime}\otimes\mathbb{1}\}$ be the two sub-algebras defined w.r.t. two different metric operators $G=\eta^{2}$ and $G^{\prime}={\eta^{\prime}}^{2}$ respectively. One defines the complementary sub-algebras $\mathcal{A}_{\pi}^{2}=\{A\in\mathcal{A}:\eta\pi_{G}(A)\eta^{-1}=\mathbb{1}\otimes\mathcal{O}_{A}\}$ and similarly, $\mathcal{A}_{\pi^{\prime}}^{2}$. This gives us two bi-partitions of the algebra $\mathcal{A}$ into $\mathcal{A}^{1}_{\pi}\vee\mathcal{A}^{2}_{\pi}$ and $\mathcal{A}^{1}_{\pi^{\prime}}\vee\mathcal{A}^{2}_{\pi^{\prime}}$.
Now we will determine the condition for the equality of the two algebras $\mathcal{A}_{\pi}^{1}\sim\mathcal{A}_{\pi^{\prime}}^{1}$. Let $X\in\mathcal{A}_{\pi}^{1}$. We know that the operators under the two Hilbert spaces are related by $\pi^{\prime}(X)=T_{U}^{-1}\pi(X)T_{U}$ such that $\eta^{\prime}T_{U}^{-1}=U\eta$. Combining these two relations we have $\eta^{\prime}\pi_{G^{\prime}}(X)\eta^{\prime-1}=U\eta\pi_{G}(X)\eta^{-1}U^{\dagger}$. Therefore $X\in\mathcal{A}_{1}^{\prime}$ if and only if $U\mathcal{O}_{X}\otimes\mathbb{1}U^{\dagger}=\mathcal{O}_{X}^{\prime}\otimes\mathbb{1}$ for the given unitary $U$. Similarly we have $\mathcal{A}_{\pi}^{2}\sim\mathcal{A}_{\pi^{\prime}}^{2}$ if and only if $U(\mathbb{1}\otimes\mathcal{O}_{Y})U^{\dagger}=\mathbb{1}\otimes\mathcal{O}_{Y}^{\prime}$ for all $Y\in\mathcal{A}^{2}_{\pi^{\prime}}$. Unitaries satisfying these conditions must always be of the form $U=U_{1}\otimes U_{2}$ (see e.g. [32, 33]). Therefore, the two partitions induced by the representations $\pi$ and $\pi^{\prime}$ are equivalent if and only if the interwining isometry $T_{U}$ satisfies the relation $\eta^{\prime}T_{U}=U_{1}\otimes U_{2}\eta$. It is straightforward to see that representations related by such an intertwiner form an equivalence class. ∎
In a typical experiment, the pseudo-Hermitian Hamiltonian is simulated through post-selection or equal gain-loss mechanisms. The form of the Hamiltonian is fixed by the experimentalist, and therefore one is restricted to a class of metric spaces compatible with this Hamiltonian. Such metric spaces are fixed by the isometry $T_{U}$, or equivalently, the unitary $U$. As a simple corollary of the above theorem, the partitioning of the system is then fixed by the choice of the metric space.
@@
Given a pseudo-Hermitian operator $H$ such that $H^{\dagger}G=GH$, the choices of the compatible representation spaces $\mathcal{H}_{G^{\prime}}$ with equivalent partitioning are those belonging to the equivalence class [G] defined by
@@
Given two cyclic representations $(\mathcal{H}_{G},\pi)$ and $(\mathcal{H}_{G^{\prime}},\pi^{\prime})$, such that $H=\pi(X)=\pi^{\prime}(X)$, therefore the representation spaces are $H$-compatible. There exists an isometry $T_{U}$ between the Hilbert spaces such that $\pi=T_{U}\pi^{\prime}T_{U}^{-1}$ which implies $[\pi(X),T_{U}]=0$. From the previous theorem, the two sub-algebra decompositions $\mathcal{A}_{\pi}^{1}\vee\mathcal{A}_{\pi}^{2}$ and $\mathcal{A}_{\pi^{\prime}}^{1}\vee\mathcal{A}_{\pi^{\prime}}^{2}$ are equivalent if and only if $\eta^{\prime}T_{U}^{-1}=(U_{1}\otimes U_{2})\eta$. ∎
@@
As a consequence of the above results, we see that while the choice of the metric does not influence the system’s properties, viz. the expectation values of the observables, it can change the properties of the subsystem, e.g. the entanglement between the subsystems. This is not unlike in the case of global unitary rotation on the system, which is simply a choice of a global reference frame, which does not change the expectation values on the system, but does change the subsystems’ properties like the entanglement between them. On the other hand, a local unitary transformation keeps the subsystem properties intact. The above theorem explains this observation from an algebraic point of view: performing a local unitary transformation in the trivial metric Hilbert space is the same as choosing a metric space from an equivalence class of metric operators, and therefore keeps the underlying algebraic partitioning the same.
@@
Tensor product structure on $\mathcal{H}_{G}$: A tensor product structure (TPS)  [17] on a Hilbert space $\mathcal{H}_{G}$ is an isomorphism $\varphi:\mathcal{H}_{G}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ where $\dim(\mathcal{H}_{1})\times\dim(\mathcal{H}_{2})=\dim(\mathcal{H}_{G})$. We know that this isomorphism given by the map  $\varphi:\,|\psi\rangle=\eta|\psi\rangle$. Consider another Hilbert space $\mathcal{H}_{G^{\prime}}$ with the metric $G^{\prime}$ and a TPS $\varphi^{\prime}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}:\psi\rightarrow\eta^{\prime}\psi$. We have also seen that there is an isometry between the two Hilbert spaces, as defined in previous sections, is $\mathcal{T}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}:\,\psi\rightarrow T_{U}|\psi\rangle$. Therefore, the map $\varphi^{\prime}\circ\mathcal{T}^{-1}:\psi\rightarrow\eta^{\prime}T_{U}^{-1}\psi$ defines another TPS over the Hilbert space $\mathcal{H}_{G}$. This gives us a family of TPS on $\mathcal{H}_{G}$ parametrized by the operator $\eta^{\prime}$, or equivalently by $U$. If the unitaries parameterizing the TPS are local transformations, i.e. of the form $U=U_{1}\otimes U_{2}$, we say they generate equivalent TPS [^2] Following the terminology of  [17]. From corollary 1.1 we see that:
@@
Remark: a family of equivalent TPS on $\mathcal{H}_{G}$ corresponds to the equivalence class of metric operators $[G]\equiv\{G^{\prime}=T_{U}^{\dagger}GT_{U};\,\eta^{\prime}T_{U}^{-1}=(U_{A}\otimes U_{B})\eta\},$ and consequently the same partitioning of the total algebra.
II.3 Partial trace
Given a genuine bi-partition $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$, we have the isomorphism $\pi(\mathcal{A}_{1})\cong\mathcal{B}(\mathcal{H}_{1})$. Therefore, the subsystem can be represented as a bounded operator on $\mathcal{H}_{1}$ in the following way: $\omega(A)|_{A\in\mathcal{A}_{1}}=tr(\rho_{\omega}\mathcal{O}_{A}\otimes\mathbb{1}_{2})=tr(\rho_{\omega}^{r}\mathcal{O}_{1}).$ The partial trace map $\rho_{\omega}\rightarrow\rho_{\omega}^{r}$ is a completely positive and trace-preserving map. From an operational point of view, the subsystem is completely defined by the algebra $\mathcal{A}_{1}$ and the value of the functional $\omega$ on this algebra. Therefore, in representations where such a genuine bi-partition is not possible, although the local description of the subsystem and operations do not exist, the subsystem is well defined. In this section, we will discuss how to define the subsystems in a non-trivial metric space and propose an extended notion of partial trace.
Earlier we had discussed the irreducibility of a set of operators. A representation of a quantum system $(\mathcal{A},\omega)$ given by the map $\pi_{\omega}$ is called an irreducible representation if the set of operators $\pi(\mathcal{A})$ is an irreducible set. Irreducibility of a representation, it turns out, depends entirely on the purity of the state $\omega$. A state is called a pure state if it can not be written as a convex combination of other states.
With these discussions, it becomes clear that the reduced state of a pure state can be a mixed state (not pure): Given a pure state $(\mathcal{A},\omega)$, the restriction of the state to a sub-algebra $\mathcal{A}_{1}$ can satisfy the condition $\omega(A)=\sum\lambda_{\alpha}\omega^{\alpha}(A),\ \forall A\in\mathcal{A}_{1}$, where $\omega^{\alpha}$ are pure states. Of course, we know that for genuine bi-partitions the partial trace of a pure state can often result in a mixed state. For a state $\omega=\sum\lambda_{\alpha}\omega_{\alpha}$, the density matrix representation under the $\pi_{\omega}$ is $\rho_{\omega}=\lambda_{\alpha}\rho_{\alpha}$, where $\rho_{\alpha}$ are pure states.
The entropy of the sub-system An important result in C-* algebra then says that a cyclic representation of $(\mathcal{A},\omega)$ is irreducible if and only if the state $\omega$ is a pure state over $\mathcal{A}$. The representation of a mixed state $(\mathcal{A},\omega)$ is therefore reducible into invariant sub-spaces $\mathcal{H}_{\omega}^{k}$ : $\pi_{\omega}=\oplus_{k}\ \pi_{\omega}^{k}$, and the Hilbert space decomposes into the invariant sub-spaces $\mathcal{H}_{\omega}=\oplus_{k}\ \mathcal{H}_{\omega}^{k}$. This decomposition can be used to construct a spectral decomposition of the state, where the invariant sub-spaces are the eigen-sub-spaces of the state: $\rho_{\omega}=\oplus e_{k}\rho_{k}$. Here $e_{k}$ are the eigenvalues of the density matrix, and if the eigen-spectrum has no degeneracy, then $\rho_{k}$ are pure states. This corresponds to the preparation of the state $\omega=\sum e_{k}\omega_{k}$. The entropy of this preparation, which we define as the Shanon entropy of the vector $\vec{e}$: $H(\vec{e})=-\sum{e_{k}\ln e_{k}}$ is therefore equal to the von-Newmann entropy of the corresponding density matrix $S_{VN}=\rho_{\omega}\ln(\rho_{\omega})$. Complications that arise with the degeneracies in the spectrum have been discussed in the appendix B (see also [34, 35, felski]).
Consider a special case, as discussed in  [37], where the subsystem of interest is the invariant subspace of the representation. Given a Hilbert space of the form $\mathcal{H}=\oplus_{\alpha}\ \mathcal{H}_{\alpha}$ with the metric $G=\oplus_{\alpha}\ G_{\alpha}$, consider a state $\rho_{\omega}=\oplus_{\alpha}\lambda_{\alpha}{\rho_{\omega}}^{\alpha}$. If all $G_{\alpha}$ operators were the same, a partial trace can be given by the density metric $\rho_{\omega}^{r}=\sum_{\alpha}\lambda_{\alpha}\rho_{\omega}^{\alpha}\,\in\mathcal{H}_{\alpha}$. However, such a summation of density matrices is not well defined when they belong to different metric spaces. In fact, one is required to perform a parallel transport of the operators to another metric space in order to add them. The isometry $T_{U}$ defined earlier provides such a metric-compatible parallel transport of operators. For this example, one can define the reduced state $\rho^{r}(\beta)=\sum_{\alpha}\lambda_{\alpha}T_{\alpha,\beta}\ {\rho_{\omega}}^{\alpha}{T_{\alpha,\beta}}^{-1}$ on the Hilbert space $\mathcal{H}_{\beta}$. Here, $T_{\alpha,\beta}$ is an isometric operation between the Hilbert spaces $\mathcal{H}_{\alpha}\to\mathcal{H}_{\beta}$ and can be written as $T_{\alpha,\beta}=\eta_{\beta}^{-1}U\eta_{\alpha}$ for a unitary operator $U$. This procedure defines a unique CPTP map between $\mathcal{B}(\oplus\mathcal{H}_{\alpha})\rightarrow\mathcal{B}(\mathcal{H}_{\beta})$ and is a consistent definition of a partial trace. Note that the reduced state thus defined is independent of the unitary operators $U$ or equivalently the choice of isometries $T_{\alpha,\beta}$. Indeed, if one chooses $U=\mathbb{1}$ for all $T_{\alpha,\beta}$, the partial trace is compatible with the partial trace by Hermitisization: $\eta_{\beta}\rho^{r}(\beta)\eta^{-1}_{\beta}=tr_{X}(\eta\rho_{\omega}\eta^{-1})$.
III Tomography and the no-signaling principle
@@
Here $\mu$ is a probability measure on the set of observables $\pi(\mathcal{A})=\mathcal{B}(\mathcal{H})$ and $\langle\mathcal{O}_{i}|\,\rho\rangle=tr(O_{i}^{\dagger}\rho)=\omega(A_{i})$ is the Hilbert-Schmidt inner product between the operators $\mathcal{O}_{i}=\pi(A_{i})$ and $\rho$. Faithfully reconstructing the state requires that the operator $W$ be full rank, which is a constraint on the measure $\mu$. It is said to be a tomographically complete measure if $W$ is full rank. Note that given a tomographically complete set of observables $\{\mathcal{O}_{i}\}$ in the Hilbert space $\mathcal{H}$, the set of observables $\{\bar{\mathcal{O}}_{i}\}$, where $\bar{\mathcal{O}_{i}}=\eta^{-1}\mathcal{O}_{i}\eta$, is a tomographically complete set in $\mathcal{H}_{G}$. Furthermore, since the norm in the two Hilbert spaces is related by $||\mathcal{O}||=||\bar{\mathcal{O}}||_{G}$, the measure in the space $\mathcal{B}(\mathcal{H}_{G})$, defined by $d\mu_{G}(\bar{\mathcal{O}})=d\mu(\mathcal{O})$, is also a probability measure. Combining these observations we can construct the sampling operator in $\mathcal{B}(\mathcal{H}_{G})$: $\overline{W}=\eta^{-1}W\eta$. Therefore, it becomes clear that if the measure $\mu$ gives a tomographically complete set of observables in $\mathcal{B}(\mathcal{H})$, then $\mu_{G}$ is tomographically complete. Furthermore, if $\mu$ is a tight frame (i.e. $W=\mathbb{1}$), then $\mu_{G}$ is also a tight frame.
@@
In an experiment, the knowledge of a state is in terms of the statistical data of the observable (e.g. $A_{i}$) and the probability of the value of the measurement being $m$, denoted by $\{m,p(m|A_{i})\}$. The connection between this data and the equation (3) is given by Born’s rule which tells us that the inner product $\langle\pi(A_{i})|\rho)\rangle$ is the expectation value of the measurement results of the observable $A_{i}$: $\sum_{m}mp(m|A_{i})=\langle\pi(A_{i})|\rho\rangle=\omega(A_{i})$. From our discussion in the previous sections, we know that this quantity is independent of the chosen cyclic representation of the system $(\mathcal{A},\omega)$ i.e. $\langle\mathcal{O}|\,\rho\rangle=\langle\bar{\mathcal{O}}|\,\bar{\rho}\rangle$.

### v3

@@ III Partitions of a system
Let
$\mathcal{A}_{\pi}^{1}$ =$=\{A\in\mathcal{A}:\eta\pi(A)\eta^{-1}=\mathcal{O}_{A}\otimes\mathbb{1}\}$ (11)
$\mathcal{A}_{\pi^{\prime}}^{1}$ =$=\{A\in\mathcal{A}:\eta^{\prime}\pi_{G^{\prime}}(A)\eta^{\prime-1}=\mathcal{O}_{A}^{\prime}\otimes\mathbb{1}\}$
be the two subalgebras of $\mathcal{A}$ defined w.r.t. two different metric operators $G=\eta^{2}$ and $G^{\prime}={\eta^{\prime}}^{2}$ respectively. One defines the complementary subalgebras $\mathcal{A}_{\pi}^{2}=\{A\in\mathcal{A}:\eta\pi_{G}(A)\eta^{-1}=\mathbb{1}\otimes\mathcal{O}_{A}\}$ and similarly, $\mathcal{A}_{\pi^{\prime}}^{2}$. This gives us two bi-partitions of the algebra $\mathcal{A}$ into $\mathcal{A}^{1}_{\pi}\vee\mathcal{A}^{2}_{\pi}$ and $\mathcal{A}^{1}_{\pi^{\prime}}\vee\mathcal{A}^{2}_{\pi^{\prime}}$.
Now we will determine the condition for the equality of the two algebras $\mathcal{A}_{\pi}^{1}\cong\mathcal{A}_{\pi^{\prime}}^{1}$. Let $X\in\mathcal{A}_{\pi}^{1}$. We know that the operators under the two Hilbert spaces are related by $\pi^{\prime}(X)=T_{U}^{-1}\pi(X)T_{U}$ such that $\eta^{\prime}T_{U}^{-1}=U\eta$. Combining these two relations we have
$\eta^{\prime}\pi_{G^{\prime}}(X)\eta^{\prime-1}=U\eta\pi_{G}(X)\eta^{-1}U^{\dagger}$ (12)
Therefore $X\in\mathcal{A}_{1}^{\prime}$ if and only if
$U\mathcal{O}_{X}\otimes\mathbb{1}U^{\dagger}=\mathcal{O}_{X}^{\prime}\otimes\mathbb{1}$ (13)
for the given unitary $U$. Similarly we have $\mathcal{A}_{\pi}^{2}\cong\mathcal{A}_{\pi^{\prime}}^{2}$ if and only if $U(\mathbb{1}\otimes\mathcal{O}_{Y})U^{\dagger}=\mathbb{1}\otimes\mathcal{O}_{Y}^{\prime}$ for all $Y\in\mathcal{A}^{2}_{\pi^{\prime}}$. Unitaries satisfying these conditions must always be of the form $U=U_{1}\otimes U_{2}$ (see e.g. [41, 42]). Therefore, the two partitions induced by the representations $\pi$ and $\pi^{\prime}$ are equivalent if and only if the intertwining isometry $T_{U}$ satisfies the relation $\eta^{\prime}T_{U}=U_{1}\otimes U_{2}\eta$. It is straightforward to see that representations related by such an intertwiner form an equivalence class. ∎
Note that taking $T_{U}=\eta^{-1}$ gives the Hermitisization of the representation in $\mathcal{H}_{G}$ and the resulting representation is in the same equivalence class. Therefore, the entanglement structure of the state $\bar{\rho}\in\mathcal{H}_{G}$ will be the same as that of the state $\rho_{\eta}=\eta\bar{\rho}\eta^{-1}\in\mathcal{H}$. We will exhibit this fact more clearly in the next section when we discuss the tomographic construction of a composite system in non-Euclidean metric space.
In a typical experiment, the pseudo-Hermitian Hamiltonian is simulated through post-selection or equal gain-loss mechanisms. The form of the Hamiltonian is therefore fixed by the experimentalist, and hence one is restricted to a class of metric spaces compatible with this Hamiltonian. Such metric spaces are fixed by the isometry $T_{U}$, or equivalently, the unitary $U$. As a simple corollary of the above theorem, the partitioning of the system is then fixed by the choice of the Hamiltonian compatible metric operator.
@@
Given a quasi-Hermitian operator $H$ such that $H^{\dagger}G=GH$, the choices of the compatible representation spaces $\mathcal{H}_{G^{\prime}}$ with equivalent partitioning are those belonging to the equivalence class [G] defined by
@@
Given two cyclic representations $(\mathcal{H}_{G},\pi)$ and $(\mathcal{H}_{G^{\prime}},\pi^{\prime})$, such that $H=\pi(X)=\pi^{\prime}(X)$, therefore the representation spaces are $H$-compatible. There exists an isometry $T_{U}$ between the Hilbert spaces such that $\pi=T_{U}\pi^{\prime}T_{U}^{-1}$ which implies $[\pi(X),T_{U}]=0$. From the previous theorem, the two subalgebra decompositions $\mathcal{A}_{\pi}^{1}\vee\mathcal{A}_{\pi}^{2}$ and $\mathcal{A}_{\pi^{\prime}}^{1}\vee\mathcal{A}_{\pi^{\prime}}^{2}$ are equivalent if and only if $\eta^{\prime}T_{U}^{-1}=(U_{1}\otimes U_{2})\eta$. ∎
@@
As a consequence of the above results, we see that while the choice of the metric does not influence the system’s properties, viz. the expectation values of the observables, it can change the properties of the subsystem, e.g. the entanglement between the subsystems. This is not unlike in the case of global unitary rotation on the system, which is simply a choice of a global reference frame, which does not change the expectation values on the system, but does change the subsystems’ properties like the entanglement between them. On the other hand, a local unitary transformation keeps the subsystem properties intact. The above theorem explains this observation from an algebraic point of view: performing a local unitary transformation in the Euclidean metric Hilbert space is the same as choosing a metric space from an equivalence class of metric operators, and therefore keeps the underlying algebraic partitioning the same.
@@
Tensor product structure on $\mathcal{H}_{G}$: A tensor product structure (TPS)  [17] on a Hilbert space $\mathcal{H}_{G}$ is an isomorphism $\varphi:\mathcal{H}_{G}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ where $\dim(\mathcal{H}_{1})\times\dim(\mathcal{H}_{2})=\dim(\mathcal{H}_{G})$. We know that this isomorphism given by the map  $\varphi:\,|\psi\rangle\to\eta|\psi\rangle$. Consider another Hilbert space $\mathcal{H}_{G^{\prime}}$ with the metric $G^{\prime}$ and a TPS $\varphi^{\prime}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}:\psi\rightarrow\eta^{\prime}\psi$. We have also seen that there is an isometry between the two Hilbert spaces, as defined in previous sections, is $\mathcal{T}:\mathcal{H}_{G^{\prime}}\rightarrow\mathcal{H}_{G}:\,\psi\rightarrow T_{U}|\psi\rangle$. Therefore, the map $\varphi^{\prime}\circ\mathcal{T}^{-1}:\psi\rightarrow\eta^{\prime}T_{U}^{-1}\psi$ defines another TPS over the Hilbert space $\mathcal{H}_{G}$. This gives us a family of TPS on $\mathcal{H}_{G}$ parametrized by the operator $\eta^{\prime}$, or equivalently by $U$. If the unitaries parameterizing the TPS are local transformations, i.e. of the form $U=U_{1}\otimes U_{2}$, we say they generate equivalent TPS [^1] Following the terminology of  [17]. From corollary 1.1 we see that:
@@
Remark: a family of equivalent TPS on $\mathcal{H}_{G}$ corresponds to the equivalence class of metric operators $[G]\equiv\{G^{\prime}=T_{U}^{\dagger}GT_{U};\,\eta^{\prime}T_{U}^{-1}=(U_{1}\otimes U_{2})\eta\},$ and consequently the same partitioning of the total algebra.
IV Tomography and the no-signalling principle
In this section, we will show that the subsystems can be constructed using usual state tomography on the system. First, we will show that the states can be constructed in $\mathcal{H}_{G}$ as well as $\mathcal{H}$ using the same tomographic data.
@@ IV Tomography and the no-signalling principle
Here $\mu$ is a probability measure on the set of observables $\pi(\mathcal{A})\cong\mathcal{B}(\mathcal{H})$ and $\langle\mathcal{O}_{i}|\,\rho\rangle=\operatorname{tr}(O_{i}^{\dagger}\rho)=\omega(A_{i})$ is the Hilbert-Schmidt inner product between the operators $\mathcal{O}_{i}=\pi(A_{i})$ and $\rho$. A faithful reconstruction of the state requires that the operator $W$ be a full-rank operator, which is a constraint on the measure $\mu$. $A$ measure that samples tomographically complete set of observables would produce a full rank sampling operator, in which case the measure is called a tomographically complete measure.
Note that given a tomographically complete set of observables $\{\mathcal{O}_{i}\}$ in the Hilbert space $\mathcal{H}$, the set of observables $\{\bar{\mathcal{O}}_{i}\}$, where $\bar{\mathcal{O}_{i}}=\eta^{-1}\mathcal{O}_{i}\eta=\pi_{G}(A_{i})$, is a tomographically complete set in $\mathcal{H}_{G}$, due to the fact that $\omega(A_{i})$ is independent of the cyclic representation. Furthermore, since the norm in the two Hilbert spaces is related by $||\mathcal{O}||=||\bar{\mathcal{O}}||_{G}$, the measure $\mu_{G}$ on the space $\mathcal{B}(\mathcal{H}_{G})$, defined by $d\mu_{G}(\bar{\mathcal{O}})=d\mu(\mathcal{O})$, is also a probability measure. Combining these observations we can construct the sampling operator in $\mathcal{B}(\mathcal{H}_{G})$:
[…]
$=\eta^{-1}W(\rho)\eta$
Therefore, if the measure $\mu$ gives a tomographically complete set of observables in $\mathcal{B}(\mathcal{H})$, then $\mu_{G}$ is also tomographically complete. Furthermore, if $\mu$ is a tight frame (i.e. if $W=\mathbb{1}$), then $\mu_{G}$ is also a tight frame. This means that given the state $\omega$ on $\{A_{i}\}\subset\mathcal{A}$, one can construct the state $\bar{\rho}$ or the state $\rho_{\eta}=\eta^{-1}\bar{\rho}\eta$ for any invertible, Hermitian operator $\eta$. In this sense, we call $\bar{\rho}$ and $\rho_{\eta}$ the tomographically equivalent pair of states.
@@
In an experiment, the knowledge of a state is given in terms of the statistical data on the observables, represented as the set of possible outcomes $\{m\}$ and the corresponding probability of the outcome when the observable $A_{i}$ is measured $\equiv\{m,p(m|A_{i})\}$. The connection between this data and the equation (15) is established by the Born’s rule which tells us that the inner product $\langle\pi(A_{i})|\rho\rangle$ is the expectation value of the measurement results of the observable $A_{i}$: $\sum_{m}mp(m|A_{i})=\langle\pi(A_{i})|\rho\rangle=\omega(A_{i})$. From our discussion in the previous sections, we know that this quantity is independent of the chosen cyclic representation of the system $(\mathcal{A},\omega)$ i.e. $\langle\pi(A)|\,\rho\rangle=\langle\pi_{G}(A)|\,\bar{\rho}\rangle_{G}=\omega(A)$.
IV.1 An example
Take the example of the spin 1/2 system. The set $\{\mathbb{1},\vec{\sigma}\}$ forms a tight frame in the Euclidean metric Hilbert space $\mathcal{H}=\mathbb{C}^{2}$ since one can write
$W(\rho)=\rho=\dfrac{1}{2}\sum_{\mu=0}^{1}\operatorname{tr}(\sigma_{\mu}\rho)\sigma_{\mu}.$ (17)
Since the Pauli matrices form the orthogonal basis for the space of observables in two dimensions, i.e. $\operatorname{tr}(\sigma_{i}\sigma_{j})=2\delta_{ij}$, they correspond to measurements of the spin in three orthogonal directions. A general observable is therefore given by $\vec{\sigma}.\hat{n}$ and is associated with a unit direction $\hat{n}$ in a 3-dimensional real vector space. Let $p(m|\hat{n}_{i})$ be the probability of getting the outcome $m$, given the spin was measured in the $\hat{n}_{i}$ direction, corresponding to the measurement of the $\sigma_{i}$ observable. By controlling the direction and magnitude of the magnetic field in the Stern-Gerlach setting (see appendix C for details of spin tomography), one produces the experimental data $\{m,p(m|\hat{n}_{i})\}$, and constructs the density matrix: $\rho=\mathbb{1}/2+\sum_{m,i}mp(m|\hat{n}_{i})\sigma_{i}$.